# Deliberate Planning in Language Models with Symbolic Representation

### v2

@@ Abstract
Planning remains a core challenge for language models (LMs), particularly in domains that require coherent multi-step action sequences grounded in external constraints. We introduce SymPlanner, a novel framework that equips LMs with structured planning capabilities by interfacing them with a symbolic environment that serves as an explicit world model. Rather than relying purely on natural language reasoning, SymPlanner grounds the planning process in a symbolic state space, where a policy model proposes actions and a symbolic environment deterministically executes and verifies their effects. To enhance exploration and improve robustness, we introduce Iterative Correction (IC), which refines previously proposed actions by leveraging feedback from the symbolic environment to eliminate invalid decisions and guide the model toward valid alternatives. Additionally, Contrastive Ranking (CR) enables fine-grained comparison of candidate plans by evaluating them jointly. We evaluate SymPlanner on PlanBench, demonstrating that it produces more coherent, diverse, and verifiable plans than pure natural language baselines. [^1] Code and data are available at https://github.com/xiongsiheng/SymPlanner.
@@ 1 Introduction
Planning is a fundamental aspect of intelligent behavior, requiring the ability to construct coherent action sequences that achieve specific goals while respecting environmental constraints. In real-world applications such as robotics, virtual agents, and interactive assistants (Jiang et al., 2024b; Jin et al., 2024; Liu et al., 2025), effective planning demands not only long-horizon decision-making but also strict adherence to logical consistency and domain-specific rules. For instance, a household robot tasked with rearranging objects must ensure that every intermediate step, such as lifting, placing, or stacking, remains valid according to physical constraints. Similarly, virtual assistants must generate interaction strategies that align with user preferences and task requirements without violating contextual limitations. These scenarios illustrate the dual challenge of maintaining both local validity (i.e., ensuring individual actions are executable) and global coherence (i.e., ensuring the sequence as a whole leads toward the intended goal). While language models (LMs) have demonstrated impressive reasoning and problem-solving capabilities, they continue to struggle in planning tasks where actions must be grounded in a verifiable and coherent world model (Valmeekam et al., 2023). Their reliance on natural language reasoning alone often leads to error-prone outputs: action sequences that appear plausible in text but fail under precise execution. This challenge becomes increasingly severe in domains that require structured, multi-step decision-making under rigid constraints, as early mistakes propagate and accumulate, resulting in incoherent or infeasible plans.
@@
Existing LM-based planning paradigms, such as chain-of-thought (CoT) reasoning (Wei et al., 2022), tree-of-thought (ToT) exploration (Yao et al., 2023), and reasoning-as-planning (RAP) (Hao et al., 2023), primarily rely on natural language to represent both states and actions. This design provides flexibility and ease of integration, as natural language serves as a universal interface for reasoning across domains. However, it also introduces several critical limitations. First, natural language descriptions are inherently ambiguous and can be interpreted in multiple ways (Jiang et al., 2024a), making it difficult for models to guarantee precise alignment between a described action and its actual effect. Second, these paradigms often lack mechanisms for validation: without grounding in an explicit environment, models cannot verify whether an action is executable, nor can they ensure that a sequence of actions consistently advances toward the goal. Third, as task complexity and planning horizons increase, small inconsistencies in language-based reasoning tend to accumulate, leading to incoherent or infeasible plans (Yang et al., 2024). For example, a model may generate a sequence that instructs stacking a block onto another that is not yet cleared, or it may repeat contradictory actions such as picking up the same object twice without putting it down. These errors are symptomatic of the gap between high-level linguistic reasoning and low-level, rule-governed execution. Consequently, while language-based approaches excel at producing diverse strategies and human-readable rationales, they often fall short in domains that demand precise, verifiable, and logically consistent multi-step planning.
@@
To address these limitations, we propose SymPlanner, a novel framework that augments language models with structured planning capabilities through integration with a symbolic environment. The symbolic environment serves as an explicit and deterministic world model, enabling actions to be executed and validated against clearly defined rules. This grounding mechanism ensures that every decision is checked for feasibility and coherence, thereby reducing the risk of cascading errors. Within SymPlanner, a policy model generates candidate actions, which are executed and verified by the symbolic world model. When an action is found invalid, the system leverages Iterative Correction (IC) to refine it using symbolic feedback, guiding the LM toward valid alternatives rather than discarding its output. Furthermore, we introduce Contrastive Ranking (CR), which enables the discriminator to evaluate multiple candidate plans jointly rather than independently, providing fine-grained comparisons that improve plan selection quality. Together, these mechanisms enhance exploration, increase robustness, and support more deliberate and reliable planning. We evaluate SymPlanner on PlanBench, a challenging benchmark for plan generation and verification, where agents manipulate blocks in structured environments. Experimental results demonstrate that SymPlanner produces plans that are significantly more valid, coherent, and diverse than natural language-only baselines. Unlike prior methods, which deteriorate quickly as task horizons grow, SymPlanner maintains competitive accuracy even for longer sequences, highlighting its scalability and robustness in multi-step reasoning tasks.
@@
  • We conduct comprehensive experiments on PlanBench, demonstrating that SymPlanner outperforms natural language-only baselines in terms of planning accuracy, coherence, and execution fidelity, particularly in long-horizon scenarios.
@@ 2 Methodology
We introduce SymPlanner, a framework that augments language models (LMs) with structured planning capabilities through integration with a symbolic environment. The framework combines three components: (i) a policy model ${\mathcal{P}}_{\pi}$ that generates candidate actions, (ii) a discriminator ${\mathcal{P}}_{\text{d}}$ that evaluates and ranks candidate trajectories, and (iii) a symbolic world model ${\mathcal{M}}_{\text{wm}}$ that deterministically simulates the effects of actions and enforces domain constraints. Together, these components allow LMs to reason in a symbolic state space, reducing ambiguity and ensuring that generated plans are both coherent and verifiable.
@@ 2.1 Problem definition
We formalize planning as the task of generating a sequence of actions $\Pi=[a_{1},a_{2},\dots,a_{T}]$ that transforms an initial state $s_{0}$ into a goal state $g$. A goal $g$ is represented as a token sequence $g=(g[1],\dots,g[L])$, and the probability of generating $g$ is given by ${\mathcal{P}}(g)=\prod_{l=1}^{L}{\mathcal{P}}(g[l]\mid g[1..l-1])$. At step $t$, $s_{t}$ denotes the symbolic state of the environment, $a_{t}$ the action chosen by the policy, and $\Pi_{t}=[a_{1},\dots,a_{t}]$ the partial plan. The symbolic world model ensures that each action $a_{t}$ is consistent with the preconditions of $s_{t-1}$ and updates the environment accordingly.
Formally, the planning process can be written as
$\Pi^{*}=\arg\max_{\Pi=[a_{1},\dots,a_{T}]}\;{\mathcal{P}}_{\pi}(\Pi\mid g,s_{0},{\mathcal{M}}_{\text{wm}})\quad\text{s.t.}\quad s_{T}=g,$
where ${\mathcal{P}}_{\pi}$ denotes the policy distribution over actions, and the constraint $s_{T}=g$ enforces that the final symbolic state must match the goal. For convenience, we define three operators as follows:
Generation.
At step $t$, the policy ${\mathcal{P}}_{\pi}$ samples $N$ candidate actions from its conditional distribution:
$\mathcal{A}_{t}=\bigl{\{}\,a_{t}^{(i)}\sim{\mathcal{P}}_{\pi}(a\mid g,s_{t-1},\Pi_{t-1})\,\bigr{\}}_{i=1}^{N}.$
Discrimination.
Given candidate actions $\mathcal{A}_{t}$ with simulated next states $\{\hat{s}_{t}^{(i)}\}$, the discriminator assigns each candidate a score $d(\cdot)$ and returns the top $b$:
$\texttt{dis}({\mathcal{P}}_{\text{d}},\mathcal{A}_{t},b)=\operatorname{TopB}\!\left(\{(a_{t}^{(i)},\hat{s}_{t}^{(i)})\}_{i=1}^{N},\;\{d(g,s_{t-1},\Pi_{t-1}\!\cup\![a_{t}^{(i)}],\hat{s}_{t}^{(i)})\}_{i=1}^{N},\;b\right).$
Symbolic simulation.
The symbolic world model ${\mathcal{M}}_{\text{wm}}$ applies action $a_{t}$ to state $s_{t-1}$, producing the next state if the action is valid, or $\bot$ otherwise:
$\hat{s}_{t}=\texttt{symSim}(g,s_{t-1},a_{t},t)=\begin{cases}T(s_{t-1},a_{t}),&\text{if }\mathrm{Valid}(s_{t-1},a_{t}),\\[4.0pt] \bot,&\text{otherwise}.\end{cases}$
@@ 2.2 System design
The core idea of SymPlanner is to structure the planning process as an iterative search in symbolic state space, guided by the generative flexibility of LMs and constrained by the deterministic rules of a world model. Rather than relying on a single roll-out, the system maintains two explicit pools: a completed pool $\mathcal{D}$, which stores plans that already reach the goal, and an unfinished pool $\mathcal{C}$, which contains partial trajectories that can be further expanded. This design allows the system to balance exploration (through unfinished trajectories) with exploitation (by recording and ranking completed plans).
@@
1:Goal $g$, initial state $s_{0}$, policy ${\mathcal{P}}_{\pi}$, world model ${\mathcal{M}}_{\text{wm}}$, discriminator ${\mathcal{P}}_{\text{d}}$, generation budget NN, step limit $T$, rollout budget $b$
@@
14:   else
@@
15:     $\mathcal{C}.\mathrm{add}((g,s_{t},\Pi_{t}))$ $\triangleright$ otherwise keep plan for further expansion
16:   end if
17:  end for
18:end for
@@
19:return $\texttt{CR}({\mathcal{P}}_{\text{d}},\mathcal{D},1)$ $\triangleright$ return best plan from completed pool
2.3 Iterative Correction
A common failure mode in LM-based planning is that the policy proposes actions that violate domain constraints, such as attempting to manipulate an object that is not present or applying an operator whose preconditions are not satisfied. In standard pipelines, such invalid actions are simply discarded, which wastes generation budget and often causes the search process to stall. The key idea of Iterative Correction is to actively repair invalid actions using feedback from the symbolic world model rather than abandoning them.
Concretely, when the simulator $\texttt{symSim}(s,a)$ returns an invalid outcome, the system provides this signal back to the policy ${\mathcal{P}}_{\pi}$. The policy then refines its prediction by conditioning on both the original context $(g,s)$ and an explicit error message produced by ${\mathcal{M}}_{\text{wm}}$ (e.g., “object not available” or “precondition not met”). This procedure is repeated up to $R$ times or until a valid action is found:
$a^{\prime}=\texttt{IC}(g,s,a),\quad\texttt{IC}:(g,s,a)\mapsto a^{\prime}\sim{\mathcal{P}}_{\pi}(a\mid g,s,\text{feedback}).$
If no valid action is obtained after $R$ retries, the branch is discarded. In this way, IC prevents early termination, reduces error propagation, and improves the stability of long-horizon planning.
2.4 Contrastive Ranking
While IC ensures that actions are valid, the discriminator must still decide which candidate trajectories are most promising. A naïve approach scores each candidate independently, but this often leads to unstable preferences and fails to capture fine-grained differences between alternatives. The key idea of Contrastive Ranking is to evaluate candidates relatively, by comparing them against one another rather than in isolation.
Formally, given a candidate set $\mathcal{G}=\{(\Pi_{i},s_{i})\}_{i=1}^{m}$, the discriminator performs pairwise comparisons to assign relative scores:
$\text{score}(i)=\sum_{j\neq i}\mathbb{I}\!\bigl{[}{\mathcal{P}}_{\text{d}}(\Pi_{i},s_{i})>{\mathcal{P}}_{\text{d}}(\Pi_{j},s_{j})\bigr{]},$
where $\mathbb{I}[\cdot]$ is an indicator function. In practice, this is implemented by computing preference logits over candidate pairs and aggregating them into a global ranking. The top-$b$ plans are then selected jointly:
@@ 2.4 Contrastive Ranking
$\texttt{CR}(\mathcal{G},b)=\operatorname{TopB}\bigl{(}\mathcal{G},\{\text{score}(i)\},b\bigr{)}.$
By comparing candidates against each other, CR produces more robust preference signals and reduces the risk of selecting suboptimal but superficially plausible trajectories. This mechanism plays a similar role to beam re-ranking in structured prediction, but with a discriminator tailored to planning contexts.

### v3

@@ Abstract
Planning remains a core challenge for large language models (LLMs), particularly in domains that require coherent multi-step action sequences grounded in external constraints. We introduce SymPlanner, a novel framework that equips LLMs with structured planning capabilities by interfacing them with a symbolic environment that serves as an explicit world model. Rather than relying purely on natural language reasoning, SymPlanner grounds the planning process in a symbolic state space, where a policy model proposes actions and a symbolic environment deterministically executes and verifies their effects. To enhance exploration and improve robustness, we introduce Iterative Correction (IC), which refines previously proposed actions by leveraging feedback from the symbolic environment to eliminate invalid decisions and guide the model toward valid alternatives. Additionally, Contrastive Ranking (CR) enables fine-grained comparison of candidate plans by evaluating them jointly. Conceptually, SymPlanner operationalizes two cognitive faculties: (i) error monitoring and repair via externalized feedback (IC) and (ii) preference formation among alternatives via pairwise comparison (CR), advancing cognitively plausible, symbol-grounded planning aligned with the rich structure in intelligent systems. We evaluate SymPlanner on PlanBench, demonstrating that it produces more coherent, diverse, and verifiable plans than pure natural language baselines [^1] Code and data are available at https://github.com/xiongsiheng/SymPlanner..
@@ 1 Introduction
Planning is a fundamental aspect of intelligent behavior, requiring the ability to construct coherent action sequences that achieve specific goals while respecting environmental constraints. In real-world applications such as robotics, virtual agents, and interactive assistants (Jiang et al., 2024b; Jin et al., 2024; Liu et al., 2025), effective planning demands not only long-horizon decision-making but also strict adherence to logical consistency and domain-specific rules. For instance, a household robot tasked with rearranging objects must ensure that every intermediate step, such as lifting, placing, or stacking, remains valid according to physical constraints. Similarly, virtual assistants must generate interaction strategies that align with user preferences and task requirements without violating contextual limitations. These scenarios illustrate the dual challenge of maintaining both local validity (i.e., ensuring individual actions are executable) and global coherence (i.e., ensuring the sequence as a whole leads toward the intended goal). While language models (LMs) have demonstrated impressive reasoning and problem-solving capabilities, they continue to struggle in planning tasks where actions must be grounded in a verifiable and coherent world model (Valmeekam et al., 2023a). Their reliance on natural language reasoning alone often leads to error-prone outputs: action sequences that appear plausible in text but fail under precise execution. This challenge becomes increasingly severe in domains that require structured, multi-step decision-making under rigid constraints, as early mistakes propagate and accumulate, resulting in incoherent or infeasible plans.
@@
Existing LLM-based planning paradigms, such as chain-of-thought (CoT) reasoning (Wei et al., 2022), tree-of-thought (ToT) exploration (Yao et al., 2023), and reasoning-as-planning (RAP) (Hao et al., 2023), primarily rely on natural language to represent both states and actions. This design provides flexibility and ease of integration, as natural language serves as a universal interface for reasoning across domains. However, it also introduces several critical limitations. First, natural language descriptions are inherently ambiguous and can be interpreted in multiple ways (Jiang et al., 2024a), making it difficult for models to guarantee precise alignment between a described action and its actual effect. Second, these paradigms often lack mechanisms for validation: without grounding in an explicit environment, models cannot verify whether an action is executable, nor can they ensure that a sequence of actions consistently advances toward the goal. Third, as task complexity and planning horizons increase, small inconsistencies in language-based reasoning tend to accumulate, leading to incoherent or infeasible plans (Yang et al., 2024). For example, a model may generate a sequence that instructs stacking a block onto another that is not yet cleared, or it may repeat contradictory actions such as picking up the same object twice without putting it down. These errors are symptomatic of the gap between high-level linguistic reasoning and low-level, rule-governed execution. Consequently, while language-based approaches excel at producing diverse strategies and human-readable rationales, they often fall short in domains that demand precise, verifiable, and logically consistent multi-step planning.
@@
To address these limitations, we propose SymPlanner, a novel framework that augments language models with structured planning capabilities through integration with a symbolic environment. The symbolic environment serves as an explicit and deterministic world model, enabling actions to be executed and validated against clearly defined rules. This grounding mechanism ensures that every decision is checked for feasibility and coherence, thereby reducing the risk of cascading errors. Within SymPlanner, a policy model generates candidate actions, which are executed and verified by the symbolic world model. When an action is found invalid, the system leverages Iterative Correction (IC) to refine it using symbolic feedback, guiding the LLM toward valid alternatives rather than discarding its output. Furthermore, we introduce Contrastive Ranking (CR), which enables the discriminator to evaluate multiple candidate plans jointly rather than independently, providing fine-grained comparisons that improve plan selection quality. Together, these mechanisms enhance exploration, increase robustness, and support more deliberate and reliable planning. We evaluate SymPlanner on PlanBench, a challenging benchmark for plan generation and verification, where agents manipulate blocks in structured environments. Experimental results demonstrate that SymPlanner produces plans that are significantly more valid, coherent, and diverse than natural language-only baselines. Unlike prior methods, which deteriorate quickly as task horizons grow, SymPlanner maintains competitive success rate even for longer sequences, highlighting its scalability and robustness in multi-step reasoning tasks.
@@
  • We conduct comprehensive experiments on PlanBench, demonstrating that SymPlanner outperforms natural language-only baselines in terms of planning success rate, coherence, and execution fidelity, particularly in long-horizon scenarios.
@@ 2 Methodology
We introduce SymPlanner, a framework that augments language models (LMs) with structured planning capabilities through integration with a symbolic environment. The framework combines three components: (i) a policy model ${\mathcal{P}}_{\pi}$ that generates candidate actions, (ii) a discriminator ${\mathcal{P}}_{\text{d}}$ that evaluates and ranks candidate trajectories, and (iii) a symbolic world model ${\mathcal{M}}_{\text{wm}}$ that deterministically simulates the effects of actions and enforces domain constraints. Together, these components allow LLMs to reason in a symbolic state space, reducing ambiguity and ensuring that generated plans are both coherent and verifiable.
@@ 2.1 Problem definition
We model a symbolic planning problem as a tuple $\mathcal{P}=(\mathcal{F},\mathcal{A},I,G)$ (Ghallab et al., 2004; Geffner & Bonet, 2013). $\mathcal{F}$ is a set of ground atoms over typed predicates; $\mathcal{A}$ is a set of grounded actions with preconditions $\mathrm{Pre}(a)\subseteq\mathcal{F}$ and add/delete effects $\mathrm{Add}(a),\mathrm{Del}(a)\subseteq\mathcal{F}$. A (symbolic) state is a set $s\subseteq\mathcal{F}$. An action $a$ is applicable in $s$ iff $\mathrm{Pre}(a)\subseteq s$, and the transition function is $\gamma(s,a)=(s\setminus\mathrm{Del}(a))\cup\mathrm{Add}(a)$. The initial state is $I$; the goal $G\subseteq\mathcal{F}$ is a partial state and is satisfied by $s$ if $s\models G$ (i.e., $G\subseteq s$).
Given a sequence (plan) $\Pi=[a_{1},\dots,a_{T}]$, we write $\gamma(I,\Pi)$ for the result of successively applying $\gamma$. Our objective is to find a plan with
$\gamma(I,\Pi)\models G,\quad\text{minimizing cost (here, plan length).}$
In SymPlanner, an LLM proposes candidate actions conditioned on the current symbolic state; a symbolic simulator serves as the world model and the arbiter of validity. Natural‑language goals $g^{\text{NL}}$ and states (if provided) are mapped to symbols by a parser $\phi_{\text{NL}\to\text{SYM}}$, yielding $G=\phi(g^{\text{NL}})$ and $I=\phi(s^{\text{NL}})$. All feasibility checks and transitions operate over symbols.
@@ 2.2 System design
The core idea of SymPlanner is to structure the planning process as an iterative search in symbolic state space, guided by the generative flexibility of LLMs and constrained by the deterministic rules of a world model. Rather than relying on a single roll-out, the system maintains two explicit pools: a completed pool $\mathcal{D}$, which stores plans that already reach the goal, and an unfinished pool $\mathcal{C}$, which contains partial trajectories that can be further expanded. This design allows the system to balance exploration (through unfinished trajectories) with exploitation (by recording and ranking completed plans).
Action proposal (generation).
Given current symbolic state $s$ and goal $G$, the policy LLM ${\mathcal{P}}_{\pi}$ proposes $N$ candidate actions
$\mathcal{A}=\{\,a^{(i)}\sim{\mathcal{P}}_{\pi}(\cdot\mid s,G,\Pi)\,\}_{i=1}^{N},$
where $\Pi$ is the partial plan used as context. Parameters are not updated at test time.
Symbolic simulation (validation & transition).
The world model ${\mathcal{M}}_{\text{wm}}$ (VAL-compatible simulator) is the arbiter of validity:
$\texttt{step}(s,a)\;=\;\begin{cases}\gamma(s,a),&\text{if }\mathrm{Pre}(a)\subseteq s,\\ \bot,\,\varepsilon,&\text{otherwise,}\end{cases}$
where $\varepsilon$ is a typed error (e.g., HandNotEmpty, NotClear(x)(x)). We prefer an explicit failure signal $\bot$ over a no-op so that invalid proposals trigger repair (IC) rather than silently stalling search.
Budgets and widths.
We distinguish two key parameters: (i) beam width $b$, the number of successors retained at each layer, and (ii) rollout budget $N$, the number of candidate actions sampled for each successor. We maintain two pools throughout the process: Open, representing the frontier of partial plans to be expanded, and Done, storing successfully completed plans.
@@
1:Goal $g$, initial state $s_{0}$, policy ${\mathcal{P}}_{\pi}$, world model ${\mathcal{M}}_{\text{wm}}$, discriminator ${\mathcal{P}}_{\text{d}}$, rollout budget $N$, step limit $T$, beam width $b$
@@
14:     $\mathcal{C}.\mathrm{pop}((g,s_{t-1},\Pi_{t}))$
15:     $b\leftarrow b-1$ $\triangleright$ reduce beam width
16:   else
@@
17:     $\mathcal{C}.\mathrm{add}((g,s_{t},\Pi_{t}))$ $\triangleright$ otherwise keep plan for further expansion
18:   end if
19:  end for
20:end for
@@
21:return $\texttt{CR}({\mathcal{P}}_{\text{d}},\mathcal{D},1)$ $\triangleright$ return best plan from completed pool
Planning process.
Why not just a classical planner?
Classical planners excel when a complete domain model and a symbolic goal are given. Our setting targets three practical gaps: (i) goals or intermediate constraints arrive in natural language and must be grounded; (ii) domain models may be incomplete, LLMs propose plausible actions while the simulator enforces safety; (iii) generalization across task variants with new objects/relations where hand‑tuning models is costly. SymPlanner exploits LLMs for proposal and uses the symbolic world model for verification, yielding a safe, language‑grounded planner.
2.3 Iterative Correction (IC)
A common failure mode in LLM-based planning is proposing actions whose preconditions are not satisfied (e.g., attempting to stack onto a non-clear block). Rather than discarding such proposals, IC uses the simulator’s typed error to repair them.
Mechanism.
Given $(s,a)$ with $\text{Sim.step}(s,a)=\bot$ and error category $\varepsilon$ (e.g., HandNotEmpty, NotClear(x)(x)), IC regenerates a corrected action conditioned on $\varepsilon$:
$a^{\prime}\sim{\mathcal{P}}_{\pi}\big(\cdot\mid s,G,\Pi,\varepsilon\big)$
repeating up to $R$ times until a valid action is found (i.e., $\text{Sim.step}(s,a^{\prime})\neq\bot$). If no valid action is obtained, the branch is pruned. Using an explicit failure signal ($\bot,\varepsilon$) instead of a silent no-op ensures repair is triggered rather than allowing the search to stall. This feedback reduces error propagation and improves long-horizon stability (Tong et al., 2024; Xu et al., 2024).
2.4 Contrastive Ranking (CR)
IC ensures candidates are valid; CR then chooses which valid successors to expand. Instead of scoring each option independently, CR compares candidates relatively via pairwise preferences, yielding more robust selection.
Mechanism.
Let $\mathcal{C}=\{(s_{i},\Pi_{i})\}_{i=1}^{m}$ be valid successors after simulation. The discriminator ${\mathcal{P}}_{\mathrm{d}}$ provides pairwise preference logits (or probabilities) for $(i\succ j)$, conditioned on the goal:
$\ell_{ij}\;=\;{\mathcal{P}}_{\mathrm{d}}\!\big((s_{i},\Pi_{i})\succ(s_{j},\Pi_{j})\,\big|\,G\big),\quad\ell_{ji}=1-\ell_{ij}.$
Aggregate a tournament-style score
$\mathrm{score}(i)\;=\;\sum_{j\neq i}\mathbb{E}\!\left[\mathbf{1}\{i\succ j\}\right]\;\approx\;\sum_{j\in\mathcal{N}_{k}(i)}\ell_{ij},$
where $\mathcal{N}_{k}(i)$ is a set of $k$ opponents sampled for efficiency, giving total complexity $O(mk)$. CR returns the $\operatorname{TopK}$ candidates by $\mathrm{score}(\cdot)$:
@@ 2.4 Contrastive Ranking (CR)
$\texttt{CR}(\mathcal{C},b)\;=\;\operatorname{TopK}\!\Bigl(\mathcal{C},\,\{\mathrm{score}(i)\}_{i=1}^{m},\,b\Bigr).$
Compared to independent scoring, pairwise contrast stabilizes preferences and reduces selection of superficially plausible but globally inferior branches (Xiong et al., 2024b).